### v2

# Learning Under Low Illumination: A Dataset and Algorithm for Traffic Sign Recognition

Abstract
Traffic signboards are vital for road safety and intelligent transportation systems, enabling navigation and autonomous driving. Yet, recognizing traffic signs at night remains underexplored due to the scarcity of realistic public datasets capturing low-light degradations and distractor classes. Existing benchmarks are predominantly daytime and do not reflect challenges such as headlight glare, motion blur, sensor noise, and vandalized or ambiguous signage. To address these gaps, we introduce INTSD, a large-scale nighttime traffic sign dataset collected across diverse regions of India. INTSD contains street-level images spanning 41 traffic signboard classes, multiple distractor categories, and varied lighting and weather conditions. The dataset is designed to support both detection and fine-grained classification under realistic nighttime scenarios. To benchmark INTSD for nighttime sign recognition, we conduct extensive evaluations using state-of-the-art detection and classification models under standardized protocols. Additionally, we present LENS-Net, a strong baseline that integrates adaptive illumination-aware detection with multimodal semantic reasoning for robust nighttime sign classification. Experiments and ablations demonstrate the challenges posed by INTSD and establish competitive baselines for future research. The dataset and code for LENS-Net is publicly available for research [^1] https://adityamishra-ml.github.io/INTSD/.
1 Introduction
Robust object detection and fine-grained classification remain core challenges in computer vision. Advances in deep learning and visual architectures have achieved remarkable progress across diverse domains [38, 40, 19, 28], supported by large-scale scene understanding datasets [26, 20, 10]. In this work, we focus on Traffic Sign Recognition (TSR), introducing a new benchmark dataset and a multimodal framework designed for robust detection and classification under challenging nighttime conditions.
Table 1: Overview of traffic sign datasets. TT100K has only 10000 images with signboard in it. §video-frames covering only 15,630 unique signs. 45 classes have more than 100 examples. The MVD data set only contains front and back signboard classes. Unique means that the same physical signboard does not appear repeatedly in multiple images. Nighttime denotes datasets primarily collected under nighttime conditions.
Dataset Images Classes Signs Attributes Region Boxes Unique Nighttime
INTSD 11016 41 14,044 tampered, blur, glare, India $\checkmark$ $\times$ $\checkmark$
advertisements,
MTSD’s attributes
CNTSSS [27] 4062 3 7687 $\times$ China $\checkmark$ $\times$ $\checkmark$
MTSD [9] 105,830 400 354,154 occluded, exterior, dummy, out-of-frame, Global $\checkmark$ $\checkmark$ $\times$
included, ambiguous
TT100K [58] 100,000 221 26,349 $\times$ China $\checkmark$ $\checkmark$ $\times$
MVD [35] 20,000 2 174,541 $\times$ Global $\checkmark$ $\checkmark$ $\times$
BDD100K [56] 100,000 1 343,777 $\times$ USA $\checkmark$ $\times$ $\times$
GTSDB [44] 900 43 852 $\times$ Germany $\checkmark$ $\times$ $\times$
RTSD [43] §179,138 156 §104,358 $\times$ Russia $\checkmark$ $\times$ $\times$
STS [21] 3777 20 5582 $\times$ Sweden $\checkmark$ $\times$ $\times$
LISA [34] 6610 47 7855 $\times$ USA $\checkmark$ $\times$ $\times$
GTSRB [45] $\times$ 43 39,210 $\times$ Germany $\times$ $\checkmark$ $\times$
BelgiumTS [31] $\times$ 108 8851 $\times$ Belgium $\times$ $\times$ $\times$
Traffic signboards play a critical role in Advanced Driver Assistance Systems (ADAS), autonomous driving, navigation, and smart city systems. However, most existing traffic sign datasets [43, 58, 9] predominantly feature daytime imagery, leaving nighttime recognition underexplored. To mitigate this gap, prior studies have used generative translation or domain adaptation techniques such as CycleGAN [57] and CyCADA [17] to synthesize nighttime scenes from daytime data. Yet, these approaches often introduce unrealistic artifacts and fail to capture complex photometric effects (such as glare, reflections, and sensor noise), resulting in a persistent domain gap that hinders robustness in real-world deployment [42, 2, 36]. Additionally, prior study shows that generative translation to nighttime conditions do not accurately assess a model’s performance during the real nighttime scenario [27].
TSR has been widely studied in computer vision research [22, 27, 11, 46]. However, these studies have largely been validated on datasets with significant limitations. These benchmarks typically lack nighttime imagery, data from diverse camera sensors, and varied weather conditions. Moreover, the traffic sign classes represented often exclude real-world adversarial examples, such as damaged or partially occluded signs. This presents a gap, underscoring the need for extensive research under more challenging and realistic operational scenarios, particularly in nighttime settings.
To address this limitation, we present INTSD, a real-world Indian Nighttime Traffic Sign Dataset. The INTSD covers a wide range of geographical area across India and features 14k+ fully annotated instances, in diverse lighting and weather conditions. It also introduces four additional classes that are frequently misclassified by TSR models but are absent in existing benchmarks. The inclusion of natural noise factors such as glare and motion blur, makes it a new standard for traffic sign detection and classification tasks.
Recent advances in Graph Convolutional Neural Networks (GCNNs) have shown significant success in image classification [51, 41] and other domain classification tasks [5, 33, 30], largely due to their inherent ability to model spatial representations. Building on this, along with the proven ability of multimodal learning to fuse complementary data sources for more robust classification [37], we also propose LENS-Net, Learned Enhancement & Nodal-Semantic Network for sign recognition. It is based on first detect and then classify principle and features: (i) an adaptive image enhancement based detector and (ii) a structured multimodal CLIP-GCNN based classifier. On INTSD, LENS-Net achieves 92.56% mAP@50 for signboard detection and macro-averaged precision of 78.89% for classification.
In summary, we make the following contributions through this research:
  1. 1.
    We present the first diverse Indian nighttime traffic sign dataset with 11k+ images across India. The dataset spans over 14K fully annotated instances, covering 41 traffic signboard and supplementary classes.
  2. 2.
    We establish extensive baselines for detection and classification on INTSD, providing future research directions.
  3. 3.
    To our knowledge, we are first [^2] Prior works have explored graph-based reasoning for traffic signs, but none have leveraged multimodal vision-language fusion or prompt-guided graph reasoning. to present a structured multimodal model, LENS-Net, that integrates vision-language representations with graph convolutional backbone for traffic signboard classification.
2 Literature Review
A wide range of traffic sign datasets support detection and recognition research. Table 1 provides overview and comparison of publicly available benchmarks. Early benchmarks such as GTSRB [45] and GTSDB [44] focused on well-lit, structured settings, while later datasets like SwedishTS [21], BelgiumTS [31], RTSD [43], and TT100K [58] expanded geographic and environmental diversity. Large-scale datasets including MVD [35], BDD100K [56], and MTSD [9] introduced complex contextual challenges. Yet, nighttime datasets remain scarce, motivating INTSD, a dedicated benchmark for low-light traffic sign detection and recognition. CNTSSS [27] recently addressed nighttime detection with a limited set of traffic sign categories. INTSD extends this direction by enabling both detection and fine-grained classification by incorporating realistic distractor categories encountered in real-world driving.
Image Enhancement: The nighttime conditions adds noise and taint the computer vision tasks. Prior studies [50] have addressed this through image enhancement techniques aimed at improving visibility before detection. Recent works [29, 4] proposed adaptive and restoration-based frameworks that jointly optimize enhancement and detection, while zero-reference mapping [14] and illumination-aware methods [46] also show notable gains in nighttime perception. Motivated by these advances, our LENS-Net detector integrates adaptive image enhancement, enabling end-to-end illumination correction tuned for sign localization.
Detection: Detection has evolved from heuristic color-shape approaches [16, 47] to deep CNN-based architectures. One-stage detectors [25, 28, 39] overcome the slowness of two-stage detectors [40, 6, 24, 13], however, cluttered nighttime signboards still challenge these models, motivating adaptive and illumination-aware detection mechanisms.
Classification: Following detection, classification remains a fine-grained visual challenge, as many signs differ only in subtle symbols or numerals (e.g., speed limit 30 vs. 80). CNN-based classifiers have long dominated this space, outperforming earlier hand-crafted pipelines [58, 15], while recent approaches use attention [1, 55] and transformer-based [12] architectures to focus on discriminative regions. Such mechanisms are critical for distinguishing visually similar signs, especially under low-light or degraded conditions. In this work, our proposed CLIP-GCNN based classifier builds upon this insight, combining cross-modal attention with graph-based reasoning for semantically coherent classification.
3 Dataset
Traffic sign distributions vary widely across regions due to differences in urban, highway, and rural infrastructure [9]. To capture this diversity, the INTSD spans six districts across India, encompassing urban and rural settings and varying environmental conditions such as weather, illumination, motion blur, and glare. Each image contains multiple sign instances, ensuring dense spatial coverage. Figure 1 illustrates representative examples and their geographic distribution. The INTSD also includes challenging scenarios in which signboards are partially occluded by foliage, obscured by other signs, partially out-of-frame, or form part of larger signboards (examples in Section 1 of supplementary material).
Figure 1: Geographical coverage and representative nighttime traffic sign examples from INTSD. The dataset spans six districts and includes both rural and urban traffic signs captured under diverse lighting and weather conditions.
3.1 Data Collection and Curation
The images are primarily extracted from videos recorded at 30fps using a low-budget ($\approx 60$ USD) action camera (AKASO Ek7000), supplemented with images from various smartphones (Samsung, IPhone, and Google Pixel), comprising a total of 11,016 images. This mix results in images of varying quality, which we organize into high-quality (smartphone) and low-quality (action camera) subsets. Of these 11,016 images, 10,004 correspond to nighttime scenes, including 6,004 images containing traffic signboards and 4,000 nighttime images without signboards. We employed a YOLOv8 model to efficiently segregate frames (sampling one frame per second) containing signs from those without signs.
Furthermore, we organize the dataset based on (i) image quality and (ii) weather conditions. Images captured using action camera tend to be of lower quality, whereas those taken with smartphones are of higher quality. These subsets facilitate image-to-image translation tasks aimed at enhancing the visual quality of traffic or driving scenes. Besides nighttime images, we provide 1,012 day images captured from the same routes in both low-quality (from action camera) and high-quality (from smartphones) formats. These images can be utilized for day-to-night or night-to-day image translation tasks. We do not include a large set of daytime images due to their availability in prior Indian dataset [52].
3.2 Traffic Sign Class Taxonomy
We follow the official traffic sign taxonomy defined by the Ministry of Road Transport and Highways (MoRTH), Government of India [32], to map signboards to 41 class labels in INTSD. This set is selected in way that ensures there is no visual overlap within classes. It is important to highlight that the proposed taxonomy is inherently flexible, allowing for the incremental expansion of INTSD.
3.3 Annotation Process
Annotating traffic signboards is a time-consuming process as it involves selecting and approving images, marking traffic signs with bounding boxes, and assigning appropriate class labels to each bounding box. To increase efficiency, we divide this process into three sequential stages: (i) image selection, (ii) manual annotation, and (iii) quality checks.
Figure 2: Examples from the INTSD. Each image highlights a representative attribute or challenge present in the dataset (e.g., glare, blur, rain, tampered signs, etc.). The icon in the bottom-right corner indicates the image source (action camera or smartphone), while the label denotes the corresponding visual attribute.
Image Selection. We employ a YOLOv8 model to initially segregate frames extracted from recorded videos into two groups: images containing traffic signboards and those without signboards. Frames are sampled at a rate of one frame per second and subsequently verified through manual inspection to ensure quality and relevance. Since the images are derived from video sequences, we discard frames with severe glare or extreme motion blur where no traffic sign is discernible. However, INTSD intentionally retains images with light-to-moderate blur or glare when traffic signs remain visible, as such conditions frequently occur in real-world nighttime driving and introduce additional challenges for TSR.
Manual Annotation. We use the VGG Image Annotator (VIA) [8] to manually draw bounding boxes around traffic signboards and assign class labels to each corresponding box. All bounding boxes are subsequently verified for accuracy and completeness. INTSD includes bounding boxes for diverse traffic sign types, including directional, mandatory, informative, rural signs, highways signs, etc.
Distractor Classes. In addition to the standard traffic sign classes, we introduce four supplementary classes to address cases that are commonly confused by the TSR models and are not represented in existing datasets: (i) the tampered class includes signboards that are broken, rusted, vandalized, or covered with pamphlets; (ii) the unknown class comprises of signboards that cannot be mapped to any existing class due to occlusion, motion blur, glare, small size, or partial visibility within the frame; (iii) the advertisement class encompasses billboards and other large advertisement boards absent from current datasets; and (iv) the warning class contains signboards indicating diversions or other road warnings. INTSD also includes images affected by moderate motion blur and light glare, reflecting real-world conditions and enhancing the model’s robustness to such visual distortions. We provide the visual examples of the distractor classes and challenges present in the INTSD in Figure 2. Refer to the Section 1 of the supplementary material for more details and visuals.
Quality Checks. The annotations in INTSD are performed following a strict and standardized set of protocols to ensure high-quality labeling. Both detection and classification annotations are triple-verified to minimize the possibility of errors. Detection errors were minimal: 3.12% false negatives were identified and added. During the second verification of class labels, 0.47% labeling error rate was observed, which dropped to 0% after the third (final) verification stage.
3.4 Data Splits
We partition the dataset into five distinct splits. The five folds contain 1188, 1226, 1175, 1189, and 1226 images with 2133, 2134, 2129, 2133, and 2128 annotated signboard instances, respectively. Stratification preserves the class distribution in each fold, preventing rare classes from being underrepresented and maintaining a consistent distribution. The same holds true for the distractor classes. The exact fold definitions and annotation files are publicly released to ensure reproducibility and facilitate standardized benchmarking.
3.5 Statistics
The INTSD exhibits several characteristics that reflect the complexity of real-world nighttime sign recognition. We analyze key dataset statistics, including sensor diversity, class distribution, image resolution, signboard scale, object density, and geographical coverage. These statistics highlight the challenges posed by the dataset and motivate the need for robust detection and recognition methods.
Sensors diversity. INTSD is captured using a mixture of action cameras and smartphones, covering images from six different sensor types. This contrasts with datasets such as TT100K [58], which rely on images captured from a single sensor setup. The resulting sensor diversity introduces variations in image quality, color response, and noise characteristics, contributing to a more diverse and realistic evaluation benchmark.
Class Distribution. Figure 3 (left) provides the class distribution of the dataset. A primary characteristic of the dataset is a significant class imbalance, following a classic long-tail distribution. A small number of classes account for a large portion of the data. For instance, the top five most frequent classes: advertisement (4,204 instances), direction (2,911), unknown (1,773), left_curve (855), and speed (851), constitute a substantial majority of all labeled signs. Conversely, rare but crucial signs such as airport, blow_horn, and cross_road are severely underrepresented. This imbalance poses a significant challenge, as standard training methods can become biased towards the majority classes, leading to poor performance on the rare but equally important minority classes. This characteristic of the data strongly motivates the need for specialized methods that can learn robust representations for both the head and tail of the distribution.
Figure 3: The treemap (on left) illustrates the class distribution. The bar graphs (on right) gives distribution of image size (golden) and signboard size (red) over $\sqrt{pixel\>area}$, and number of signboards per image (green).
Image Sizes. Figure 3 (golden/right) depicts the bimodal distribution of the image sizes in the INTSD. It emphasizes that the images primarily come in two distinct size ranges: a large number of smaller images (around $\sqrt{area}\approx 1100$) and another group of much larger images (around $\sqrt{area}\approx 3700$).
Signboard Sizes. Figure 3 (red/right) illustrates the long-tail distribution of signboard sizes in INTSD. As evident, the majority of signboards are relatively small, with a progressively smaller proportion of larger signboards. This inherent size imbalance poses a significant challenge for detection models, as small signboards contain limited visual information and are more susceptible to being overlooked by standard detection architectures. Such a distribution highlights the importance of employing strategies like scale-aware feature extraction and data augmentation to mitigate the bias towards more prevalent size categories.
Signboards per Image. Figure 3 (green/right) also depicts a long-tail distribution of object density across the dataset with average of 2.32 signboards per image ($14044/6004$). Most images contain only a few signboards (1-4), whereas images with a higher number of signboards, corresponding to more cluttered scenes, are relatively less. This imbalance presents additional challenges for detection models, as densely populated scenes can introduce occlusions and complex spatial arrangements, making accurate localization and classification more difficult. Addressing this variation in object density requires models capable of robustly handling both sparse and cluttered scenarios.
Geographical Distribution. Figure 1 illustrates the geographical distribution of the images in our dataset. The images span six districts and encompass both rural and urban settings, as well as a diverse range of lighting conditions, including well-lit and poorly-lit environments. This diversity ensures that the dataset captures real-world variability.
Overall, these statistics demonstrate that INTSD captures diverse real-world conditions, making it a challenging benchmark for nighttime sign recognition.
Figure 4: Overview of the proposed LENS-Net architecture. The detector (top) integrates adaptive image enhancement for joint illumination correction and sign localization. The classifier (bottom) employs a structured multimodal CLIP-GCNN module that leverages cross attention and graph-based reasoning for sign classification.
4 Method
The proposed method, LENS-Net, comprises of a two-stage architecture designed for robust signboard detection and subsequent fine-grained classification:
  • An adaptive pre-processing LENS-Net detector ($\mathcal{D}$) identifies object bounding boxes from a raw image.
  • A multimodal graph LENS-Net classifier ($\mathcal{C}$), classifies the content of each cropped bounding box.
This decoupled design allows each stage to be optimized for its specific task: $\mathcal{D}$ for robust localization under varying imaging conditions, and $\mathcal{C}$ for high-accuracy classification using deep semantic and attribute-based features. Figure 4 provides an overview of the proposed architecture.
4.1 LENS-Net Detector
YOLO models demonstrate strong capability in detecting small and densely packed objects, owing to advances in multi-scale feature aggregation and attention-based refinement mechanisms that preserve fine-grained spatial details. Motivated by these strengths, we adopt YOLOv8 as our signboard detector. However, the INTSD comprises images captured under highly varying illumination conditions. While some scenes exhibit good lighting due to street lamps and vehicle headlights, others are severely underexposed in the absence of such sources. This variability poses a significant challenge for a static data pre-processing pipeline. To address this issue, we wrap the detector’s backbone with a PreprocessWrapper that applies a lightweight parameter prediction head $\Phi_{\text{head}}$ that estimates per-image enhancement parameters (gamma, brightness, unsharp masking). These parameters are applied via differentiable transforms, producing preprocessed images which are then fed for detection. We refer the interested readers to the supplementary material (Section 3) for more details on $\Phi_{\text{head}}$ and adaptive image enhancement.
To ensure that preprocessing module learns meaningful transformations and avoids collapsing to trivial solutions, we employ two key strategies. (i) Stochastic Augmentation: During training, we apply a diverse set of strong, non-differentiable augmentations (photometric jitter, JPEG compression, Gaussian noise, blur, and small rotations) before the image is passed to the PreprocessWrapper. It forces $\Phi_{\text{head}}$ to learn to invert these degradations and normalize the image to a consistent domain for the detector. (ii) Preprocessing Loss: We apply an L2 regularization based loss ($\mathcal{L}_{\text{preproc}}$) that penalizes the predicted scaled parameters for deviating from their (no operation) default values: $\mathcal{L}_{\text{preproc}}\leftarrow\mathbb{E}\left[\sum_{i\in\{\gamma,\alpha,\zeta\}}\lambda_{i}(p_{i}-p_{\text{default},i})^{2}\right]$, where $\lambda_{i}$ ($\lambda_{i}\in(0,0.01)$) are per-parameter loss weights. This loss is added to the standard YOLOv8 detection loss, $\mathcal{L}_{\text{detect}}$, during training. Therefore, the net loss for detector is: $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{detect}}+\mathcal{L}_{\text{preproc}}$.
LENS-Net detector is trained using a zero-one scheduler: (i) head-only optimization with frozen backbone, and (ii) joint end-to-end training. The detector’s bounding-box output ($\mathbf{I}_{\text{crop}}$) is forwarded to the classifier for recognition.
4.2 LENS-Net Classifier
After the detector localizes a traffic sign and outputs the corresponding cropped region, the classifier is tasked with accurately identifying the specific category. We design a tri-modal fusion classifier that combines: (i) robust, global semantic features from a frozen CLIP vision encoder, (ii) prompt-conditioned textual priors from CLIP text encoder, and (iii) structured attribute reasoning using a small graph. The complete architecture is detailed in Algorithm 1. The model processes the $\mathbf{I}_{\text{crop}}$ in two parallel branches that are fused for a final prediction.
Branch 1: Text-Image Fusion. We process $\mathbf{I}_{\text{crop}}$ to obtain the corresponding global image embedding $\mathbf{v}$, which is fused with semantic context derived from a set of pre-defined textual prompts $\mathbf{P}$. These prompts are encoded into embeddings $\mathbf{T}\in\mathbb{R}^{n(\mathbf{P})\times D}$ using the frozen CLIP text encoder (Algorithm 1, line 2-3). $\mathbf{v}$ and $\mathbf{T}$ are fused with cross-attention (Equation 1). This produces a prompt-conditioned image feature CA1, a refined global vector that already blends visual information with textual priors. CA1 is processed with residual connection (Res), Layer Normalization (LN), and position-wise feed-forward network (FFN) to produce the text-conditioned feature $\mathbf{f}_{\text{CA1}}$ (Equation 2).
Algorithm 1 LENS-Net Classifier
1:Image crop $\mathbf{I}_{\text{crop}}$ (from LENS-Net Detector) 2:Frozen $E^{img}_{clip},\;E^{text}_{clip}$; Static text prompts $\mathbf{P}$ $\triangleright$ CLIP encoders 3:Class logits $(\mathbf{I}_{\text{crop}})_{\text{class}}$ 4:procedure ClassifierForward($\mathbf{I}_{\text{crop}}$) Get Foundational Embeddings 5:  $\mathbf{v}\leftarrow\text{L2Normalize}(E^{img}_{clip}(\mathbf{I}_{\text{crop}}))$ $\triangleright$ Image vector 6:  $\mathbf{T}\leftarrow\text{GetPrecomputedTextEmbeddings}(\mathbf{P})$ $\triangleright$ Prompt matrix Branch 1: Text-Image Fusion 7:  $\mathbf{f}_{\text{CA1}}\leftarrow\text{TextImageFusionModule}(\mathbf{v},\mathbf{T})$ $\triangleright$ Equation 2 Branch 2: Graph Attribute Module 8:  $\mathbf{H}^{(L)}\leftarrow\text{GraphAttributeModule}(\mathbf{v})$ $\triangleright$ Handles node creation & GCNN Final Fusion & Classification 9:  $\mathbf{f}_{\text{final}}\leftarrow\text{FinalFusionModule}(\mathbf{f}_{\text{CA1}},\mathbf{H}^{(L)})$ $\triangleright$ Equation 6 10:  $(\mathbf{I}_{\text{crop}})_{\text{class}}\leftarrow\text{ClassifierHead}(\mathbf{f}_{\text{final}})$ 11:  return $(\mathbf{I}_{crop})_{\text{class}}$ $\triangleright$ Output class category 12:end procedure 13:function GraphAttributeModule($\mathbf{v}$) 14:  $\mathbf{n}_{G}\leftarrow\mathbf{v}$ 15:  $\mathbf{n}_{S}\leftarrow(\text{Softmax}(\text{MLP}_{S}(\mathbf{v})))\cdot\mathbf{E}_{S}$ $\triangleright$ $\mathbf{E}_{S}$: learnable shape embedding matrix 16:  $\mathbf{n}_{C}\leftarrow(\text{Softmax}(\text{MLP}_{C}(\mathbf{v})))\cdot\mathbf{E}_{C}$ $\triangleright$ $\mathbf{E}_{C}$: learnable color embedding matrix 17:  $\mathbf{H}^{(0)}\leftarrow\text{Stack}(\mathbf{n}_{G},\mathbf{n}_{S},\mathbf{n}_{C})$ $\triangleright$ Graph construction 18:  $\mathbf{H}^{(L)}\leftarrow\text{SimpleGCNN}(\mathbf{H}^{(0)})$ $\triangleright$ Updated node features 19:  return $\mathbf{H}^{(L)}$ 20:end function
$\mathbf{CA1}=\text{MHA}_{\text{CA1}}(\mathbf{Q}=\mathbf{v},\mathbf{K}=\mathbf{T},\mathbf{V}=\mathbf{T})$ (1)
$\mathbf{f}_{\text{CA1}}=\text{FFN}(\text{LN}(\mathbf{v}+\mathbf{CA1})),\>\>\mathbf{f}_{\text{CA1}}\in\mathbb{R}^{D}$ (2)
Equation 1 and 2 represents the TextImageFusionModule in Algorithm 1 (line 4). $\mathbf{f}_{\text{CA1}}$ aligns the image’s representation with textual distinctions that are useful for traffic-sign recognition (e.g., color/shape descriptions, presence of numbers or arrows, etc.).
Branch 2: Graph Attribute Module. This branch explicitly models the two most salient attributes of signs: shape and color. As shown in the GraphAttributeModule (Algorithm 1, lines 10-17), from the global embedding $\mathbf{v}$, we predict soft distributions over learnable shape ($\mathbf{E}_{S}$) and color ($\mathbf{E}_{C}$) embedding tables. It turns the predicted distributions into continuous node vectors, i.e., instead of making a hard, non-differentiable argmax decision (e.g., this is 100% a circle), we represent uncertainty (e.g., [0.85 (circle), 0.1 (octagon), 0.05 (other)]).
For fine-grained sign classification, we construct three nodes: global ($\mathbf{n}_{G}$), shape ($\mathbf{n}_{S}$), and color ($\mathbf{n}_{C}$). Additional details on node construction are provided in Section 3 of the supplementary material. These form a structured 3-node graph: $\mathbf{H}^{(0)}=[\mathbf{n}_{G},\mathbf{n}_{S},\mathbf{n}_{C}]^{T}$, refined using a 3-layer GCNN (SimpleGCNN on line 15 in Algorithm 1) operating on node matrix $\mathbf{N}$ $(\mathbf{N}=[\mathbf{n}_{G},\mathbf{n}_{S},\mathbf{n}_{C}],\mathbf{N}\in\mathbb{R}^{D\times 3})$. Let $\mathbf{A}\in\mathbb{R}^{3\times 3}$ be the fully-connected and symmetrically normalized adjacency (Equation 34). The GCNN outputs updated node features $\mathbf{H}^{\text{L}}$.
$\hat{A}=D^{-\frac{1}{2}}AD^{-\frac{1}{2}},\qquad D_{ii}=\sum_{j=1}^{3}A_{ij}.$ (3)
$\mathbf{H}^{(l+1)}=\hat{A}\,\big(\mathbf{H}^{(l)}W^{(l)}+b^{(l)}\big),\qquad\mathbf{H}^{(0)}=\mathbf{N}^{T},$ (4)
where $\mathbf{H}^{(l)}\in\mathbb{R}^{3\times D}$ is node features at layer $l$, and $W^{(l)},b^{(l)}$ are the layer parameters. After $L$ layers, the GCNN outputs node features: $\mathbf{H}^{(L)}\in\mathbb{R}^{3\times D}.$
Final Fusion. The text-conditioned feature $\mathbf{f}_{\text{CA1}}$ queries the GCNN-refined node features $\mathbf{H}^{(L)}$ via a second cross-attention module (Equation 5; Algorithm 1, line 6). CA1 lets textual priors refine the image and CA2 lets the refined image go back and attend to structured attribute features. The output is processed with residual normalization and FFN to obtain the final feature $\mathbf{f}_{\text{final}}$ (Equation 6). The final vector $\mathbf{f}_{\text{final}}$ is passed to an MLP head (Algorithm 1, line 7) and produces logits ($(\mathbf{I}_{\text{crop}})_{\text{class}}\in\mathbb{R}^{N}$) over the target traffic-sign classes.
$\mathbf{CA2}=\text{MHA}_{\text{CA2}}(\mathbf{Q}=\mathbf{f}_{\text{CA1}},\mathbf{K}=\mathbf{H}^{\text{L}},\mathbf{V}=\mathbf{H}^{\text{L}})$ (5)
$\mathbf{f}_{\text{final}}=\text{FFN}(\text{LN}(\mathbf{f}_{\text{CA1}}+\mathbf{f}_{\text{CA2}}))$ (6)
Classifier Training Objective. The classifier is trained using class-balanced focal loss [25] (Equation 7) to mitigate long-tailed class imbalance in INTSD. Class weights $\alpha_{t}$ are set proportional to the inverse square root of class frequency.
$\mathcal{L}_{\text{FL}}(p_{t})=-\alpha_{t}(1-p_{t})^{\gamma}\log(p_{t})$ (7)
where $p_{t}$ is the model’s predicted probability for the ground-truth class, and $\gamma$ is the focusing parameter. To further mitigate imbalance, we employ a rarity-based augmentation. Implementation details and augmentation strategies are mentioned in the supplementary material (Section 2).
5 Results
Experimental Settings. We adopt a stratified 5-fold cross-validation strategy for training and evaluation, thereby reducing bias in the evaluation process. We choose mAP@50 and mAP@50:95 as detection evaluation metrics and macro-averaged precision (A.P.), macro-averaged recall (A.R.), and accuracy (Acc.) as classification evaluation metrics. We also report frames per second (FPS) to evaluate real-time performance and deployment feasibility. For fair comparison, FPS is measured end-to-end, from raw image loading to final predictions, including preprocessing, forward pass, post-processing, and CUDA synchronization. Each FPS value is averaged over five independent runs.
Baseline Results. In Tables 2 and 3, we benchmark INTSD using a diverse set of strong baselines covering multiple architectural paradigms. For detection, we evaluate: (i) two-stage detectors (Faster R-CNN [40]), (ii) one-stage detectors (YOLOv8 [18], YOLOv10-L [53], YOLOv12-L [48], GOLD-YOLO-L [54]), (iii) low-light–oriented detection methods (DAI-Net [7], YOLO-LLTS [27], YOLO-TS [3]), and (iv) TSR approaches (YOLO-LLTS [27], YOLO-TS [3]).
For classification, we compare against: (i) convolutional and transformer-based models (DeIT [49]), (ii) vision-language models (CLIP [37], BLIP [23]), (iii) YOLO based methods (YOLOv10-L [53], YOLOv12-L [48]) and (iv) task-specific TSR approaches (Rishabh et al. [52], YOLO-TS [3]). All models are trained and evaluated under identical data splits and reporting protocols.
Detection. Table 2 reports detection performance on INTSD across two-stage, one-stage, and low-light–oriented detectors. Modern YOLO variants achieve strong results, reflecting the dataset’s suitability for contemporary detection architectures. LENS-Net attains the highest mAP@50 (92.56%) and mAP@50:95 (79.54%), while maintaining competitive inference speed. The consistent performance gap over both generic and low-light–specific detectors highlights the challenges posed by diverse nighttime conditions and validates the benchmark’s ability to differentiate strong methods.
Table 2: Detection performance on the INTSD.
Model mAP@50 (%) mAP@50:95 (%) FPS Parameters (M)
Total Trainable
Faster R-CNN [40] 62.60 43.22 23.38 41.50 41.50
DAI-Net [7] 87.18 62.56 20.15 49.83 49.83
YOLOv12-L [48] 82.84 64.93 59.08 26.48 26.48
GOLD-YOLO-L [54] 88.70 72.11 42.64 74.99 74.99
YOLO-LLTS [27] 90.00 73.65 56.34 9.87 9.87
YOLOv10-L [53] 90.00 75.81 64.39 25.83 25.83
YOLOv8 [18] 90.13 77.21 83.71 11.14 11.14
YOLO-TS [3] 90.82 77.27 62.52 23.39 23.39
LENS-Net 92.56 79.54 82.27 11.22 11.22
Table 3: Classification results on the INTSD.
Model A.P. (%) A.R. (%) Acc. (%) FPS Parameters (M)
Total Trainable
YOLOv10-L [53] 37.24 26.51 70.32 64.37 25.83 25.83
YOLOv12-L [48] 45.70 39.60 67.66 59.14 26.48 26.48
BLIP [23] 49.33 47.81 69.90 75.37 4060.87 8.93
CLIP [37] 65.25 68.11 82.49 72.37 434.04 12.85
DeIT [49] 65.89 59.69 82.13 136.98 85.83 85.83
Rishabh et al. [52] 74.49 69.98 88.21 60.25 26.87 26.85
YOLO-TS [3] 75.08 74.43 89.70 63.36 23.39 23.39
LENS-Net 78.89 75.56 88.59 62.44 439.37 11.75
Classification. Table 3 presents classification performance across convolutional, transformer-based, vision-language, and task-specific TSR models. Vision-language approaches such as CLIP and BLIP provide strong semantic baselines, while task-specific models (e.g., Rishabh et al. [52], YOLO-TS [3]) achieve higher task-adapted accuracy. LENS-Net achieves the best macro-averaged precision (78.89%) and recall (75.56%), demonstrating the effectiveness of structured multimodal reasoning under long-tailed and low-illumination conditions. Ablation study (see supplementary material, Section 4) confirms the contribution of key components in LENS-Net classifier. These results establish competitive reference baselines for future research on INTSD. Figure 5 gives the predictions of LENS-Net on INTSD along with false positives and false negatives.
Figure 5: Qualitative results of LENS-Net on INTSD. The green, red, and blue boxes indicate true positives, false negatives, and false positives, respectively. The numbers on the boxes represent the corresponding class IDs.
Generalization to CNTSSS. Table 4 reports cross-dataset evaluation on CNTSSS [27]. While several methods achieve strong performance due to its limited class taxonomy, INTSD produces larger performance variation across models, reflecting increased semantic diversity and intra-class ambiguity. Notably, LENS-Net achieves competitive detection performance and the highest classification accuracy, demonstrating robustness across differing nighttime benchmarks.
Table 4: Detection and classification performance on CNTSSS. Rishabh et al. and YOLO-LLTS are classification-only and detection-only models, respectively.
Model Detection (%) Classification (%)
mAP@50 mAP@50:95 FPS Parameters (M) A.P. A.R. Acc. FPS Parameters (M)
YOLOv10-L 68.84 46.56 100.51 26.77 84.19 84.15 86.02 100.32 26.77
Rishabh et al. 90.00 89.33 91.68 63.52 26.85
YOLO-LLTS 76.30 54.15 72.17 9.85
YOLO-TS 76.47 54.21 102.51 23.39 83.92 84.08 86.46 102.54 23.39
LENS-Net 69.17 46.62 151.75 11.13 90.67 90.02 92.10 70.60 11.73
6 Conclusion
We present and benchmark INTSD, the first diverse Indian nighttime dataset for traffic sign detection and classification, and LENS-Net, a multimodal framework that unifies adaptive enhancement, vision-language priors, and graph-based reasoning. Experimental results demonstrate LENS-Net’s superior performance, providing a valuable benchmark and foundation for future research on robust perception in adverse illumination environments.